# COCOLA: Coherence-Oriented Contrastive Learning of Musical Audio Representations

### v2

@@ Abstract
We present COCOLA (Coherence-Oriented Contrastive Learning for Audio), a contrastive learning method for musical audio representations that captures the harmonic and rhythmic coherence between samples. Our method operates at the level of stems (or their combinations) composing music tracks and allows the objective evaluation of compositional models for music in the task of accompaniment generation. We also introduce a new baseline for compositional music generation called CompoNet, based on ControlNet, generalizing the tasks of MSDM, and quantify it against the latter using COCOLA. We release all models trained on public datasets containing separate stems (MUSDB18-HQ, MoisesDB, Slakh2100, and CocoChorales).
1 Introduction
Recently, there have been significant advances in music generation in the continuous domain [1, 2, 3, 4, 5], thanks to the impressive development of generative models [6, 7, 8]. In addition to producing high-quality tracks of increasing length [5], these models offer precise semantic control through textual conditioning [9, 10]. However, they are limited as tools for musical composition, since they output a final mix containing all stems. To overcome this, a new range of compositional generative models is emerging [11, 12, 13], where (i) the generative tasks are defined at the stem level and (ii) their usage is iterative/interactive. The most important application of these models is accompaniment generation, where, given multiple conditioning sources (combined or not), the model is asked to output a new set (or a mixture) of coherent stems. Although previous models could generate accompaniments [14, 15], they could not be used iteratively (acting sequentially on the model’s outputs) in a composition process.
@@ 1 Introduction
Figure 1: Illustration of COCOLA score. COCOLA is a contrastive model able to estimate the coherence between instrumental tracks and generated accompaniments.
To this end, we propose a novel contrastive model called COCOLA (Coherence-Oriented Contrastive Learning for Audio), which can evaluate the coherence between conditioning tracks and generated accompaniments (Figure 1). The model is trained by maximizing the agreement between disjoint sub-components of an audio window (sub-mixtures of stems) and minimizing it on sub-components belonging to different windows. With the model, we define a COCOLA score as the similarity between conditioning tracks and accompaniments in the embedding space.
Additionally, given the scarcity of open-source compositional music models (to our knowledge, only MSDM is available publicly [11]), we introduce and release [^1] https://github.com/gladia-research-group/cocola an improved latent diffusion model called CompoNet, based on ControlNet [17], which generalizes the tasks of previous compositional models. We benchmark CompoNet against MSDM with the COCOLA score and FAD [16] on accompaniment generation, showcasing better performance.
After discussing related work in Section 2, we introduce COCOLA and CompoNet in Section 3. We describe the experimental setup in Section 4 and present the results in Section 5. We conclude the article in Section 6.
Figure 2: The COCOLA training procedure (single stem case). We first randomly crop windows of size $L$ from a batch of $K$ tracks (depicted on the left). As a second step, we randomly select two distinct stems in each window. For example, in the first window we select $\mathbf{x}^{1}_{1}$ (Guitar) and $\mathbf{x}^{1}_{3}$ (Drums). Thus, we embed all selected stems with the COCOLA encoder $f_{\theta}$, obtaining latent representations. For example, we obtain $\mathbf{h}^{1}_{1}$ and $\mathbf{h}^{1}_{2}$ from the first window. Finally, we compute the contrastive loss (Eq. (3)) considering embedings belonging to the same window as positive pairs and combinations of embeddings between different windows as negative pairs.
2 Related Work
2.1 Contrastive Methods for Audio
@@
Supervised contrastive learning methods are typically cross-modal, requiring labeled information alongside audio data. In early works, the labeled information was in the form of simple tags, while the loss used to align embeddings of audio segments and tags was the triplet loss [20]. Within the same data setting,[21] used the contrastive loss of SimCLR [22]. With the advent of the transformer architecture [23], using complex sentences instead of simple tags became feasible. MuLaP [24] is the first model to train a common representation between audio and sentences in the musical domain. In such work, the audio and text are processed by a joint transformer encoder, conveying information about the two modalities through cross-attention layers. Although it is not a contrastive model per se, an audio-text matching loss uses negative examples to encourage the model to focus on aligned pairs. More recent works [25, 26, 9, 27], consider separate textual and audio encoders, which makes it possible to use the two branches independently at inference time.
@@
Self-supervised representation learning methods [28, 29, 30, 31] build embedding spaces targeting structural information extracted from the audio data itself. In [32], the authors build positive examples for a triplet loss by augmenting with Gaussian noise, time and frequency translations, and sampling with time proximity. They also consider example mixing. While we compare coherent mixes in our method (Section 3.1), in [32], positive pairs are not coherence-related (e.g., mixing siren and dog sounds). As in the supervised case, following [22], multi-class cross-entropy losses are employed [33, 34, 35]. In COLA [33], the authors train an embedding model with contrastive loss using the simple criterion of sampling positive pairs only from the same audio track (still employing Gaussian noise), outperforming a fully supervised baseline in a plethora of tasks. [36] pairs mixtures with sources extracted via source separation.
The proposed COCOLA method shares aspects of both supervised and self-supervised approaches. Given that stems are pre-separated, we cannot consider the method purely self-supervised. At the same time, we process such data with a uni-modal encoder, as is the case of self-supervised methods.
2.2 Compositional Waveform Music Generation
Compositional music generation in the waveform domain (as opposed to symbolic domains such as MIDI or sheet music [37]) was introduced by [11], proposing the Multi-Source Diffusion Model (MSDM). Such a model captures the joint distribution of a fixed set of sources (e.g., Bass, Drums, Guitar, and Piano) using a (score-based) diffusion model [38, 39, 7, 6, 40]. At inference time, it is possible to perform unconditional generation of all sources, accompaniment generation of a subset given the complement, and source separation.
Following this initial work, AUDIT [41] proposes a diffusion model conditioned by a T5 Encoder [10], trained with instructions that allow the addition, removal (drop), and replacement of sources in an input audio mixture. This model operates on general audio signals with weak dependencies between the sources (e.g., environmental sounds). While MSDM is an unconditional generative model that processes single sources in parallel, AUDIT is a conditional generative model that processes mixtures sequentially.
InstructME [12] introduces AUDIT in the musical setting of MSDM, where sources are highly interdependent. Besides conditioning on chords, the fundamental difference lies in how the audio input is provided to the model: while in AUDIT, the input and output are two channels of the tensor that the diffusion model processes (the input is inpainted at inference time, similarly to how accompaniment generation is performed in MSDM), in InstructME, the input is processed by a convolutional network that follows the structure of the diffusion model U-Net [42] encoder. The features processed by the latter are aggregated with the features of the U-Net encoder (authors do not specify if they sum or inject the features).
StemGen [13] generates a single accompaniment source, given an instrument tag and an input audio mixture, via a masked music language model [4].
@@ 2.2 Compositional Waveform Music Generation
Another line of research focuses on inference-only methods for compositional music tasks, given pre-trained generative models. Based on generative source separation via Bayesian inference [43, 44, 45], Generalized Multi-Source Diffusion Inference (GMSDI) [46] performs the tasks of MSDM, requiring models trained only with mixtures and text, by separating sources while generating them.
Our proposed CompoNet baseline (Section 3.3) is a sequential conditional model that, unlike previous models, can perform all compositional music generation tasks (Figure 3). Differently than InstructME, the model is conditioned via a ControlNet adapter [17], which enables fine-tuning of diffusion models pre-trained with a large quantity of mixtures. Table 1 summarizes a comparison between CompoNet and current music compositional models.
3 Method
3.1 Stem-Level Contrastive Learning
@@
$\mathbf{m}^{k}_{1}=\sum_{\mathbf{x}_{n}^{k}\in X_{1}^{k}}\mathbf{x}^{k}_{n},\qquad\mathbf{m}^{k}_{2}=\sum_{\mathbf{x}_{n}^{k}\in X_{2}^{k}}\mathbf{x}^{k}_{n}$ (1)
When $X_{1}^{k},X_{2}^{k}$ are singletons, the sub-mixes are simply two stems in the window (single stem case). We work with sub-mixes because current compositional music generation methods [13] operate over them, including our proposed CompoNet (see Section 3.3). Like in COLA[33], we use a convolutional audio-only encoder [^2] In our notation, we incorporate into $f_{\theta}$ any domain transform preceding or following the convolutional network operations, like the (pre) mel-filterbank map and the (post) projection head $g$ in COLA. $f_{\theta}:\mathbb{R}^{L}\to\mathbb{R}^{d}$, mapping $\mathbf{m}^{k}_{1}$ and $\mathbf{m}^{k}_{2}$ to lower-dimensional embedding vectors $\mathbf{h}^{k}_{1}=f_{\theta}(\mathbf{m}^{k}_{1})$ and $\mathbf{h}^{k}_{2}=f_{\theta}(\mathbf{m}^{k}_{2})$, with $d$ the embedding dimension.
The COCOLA training procedure maximizes the agreement between pairs $\mathbf{h}^{k}_{1},\mathbf{h}^{k}_{2}$ of sub-mixes embeddings in the same window. It decreases it for pairs $\mathbf{h}^{k}_{1},\mathbf{h}^{j}_{2}$ ($j\neq k$) of sub-mixes embeddings in different windows. As in COLA, we use a bilinear similarity metric:
@@
Figure 3: Inter-stem compositional generation tasks. $Y$ and $X$ represent the input and output stem sub-sets, respectively.
Model Task Methodology Input Output Coherence
UG AG SS EA ER
MSDM [11] ✓(MS) ✓(MS) Training / Supervised Multi / Source Multi / Source
GMSDI [46] ✓(MS) ✓(MS) Inference / Weakly Supervised Multi / Sub-mix Multi / Sub-mix
StemGen [13] ✓(1S) Training / Supervised Single / Sub-mix Single / Source
Audit [41] ✓(Remove 1S) ✓(1S) ✓(1S) Training / Supervised Single / Sub-mix Single / Sub-mix
InstructME [12] ✓(Extract 1S; Remove 1S) ✓(1S) ✓(1S) Training / Supervised Single / Sub-mix Single / Sub-mix
@@
Table 1: Compositional audio models comparison. The various tasks are illustrated in Figure 3. 1S vs MS: the task operates on one vs multiple sources at a time. Multi vs Single on Input / Output: the model accepts multiple vs single inputs / outputs. Source vs Sub-mix on Input / Output: the model processes single sources or sub-mixes as inputs / outputs.
3.2 COCOLA Score
Equipped with the encoder $f_{\theta}$, we can quantify the coherence of the accompaniments generated by a generative model $p_{\phi}(\mathbf{x}\mid\mathbf{y})$, where $\mathbf{y}$ is the conditioning variable (the input) and $\mathbf{x}$ is the modeled variable (the output). The model’s variables can be either a set of stems or sub-mixes. Given the input $\mathbf{y}$, the model $p_{\phi}$ generates an output $\tilde{\mathbf{x}}\sim p_{\phi}(\mathbf{x}\mid\mathbf{y})$. We can compute the coherence between $\mathbf{y}$ and $\tilde{\mathbf{x}}$ by first embedding the two vectors $\mathbf{h}_{\mathbf{y}}=f_{\theta}(\mathbf{y})$ and $\mathbf{h}_{\tilde{\mathbf{x}}}=f_{\theta}(\tilde{\mathbf{x}})$ (summing the stems beforehand if considering a set of stems). We define the COCOLA score (CCS) between $\mathbf{x}$ and $\tilde{\mathbf{y}}$ as:
$\text{CCS}(\mathbf{y},\tilde{\mathbf{x}})=\text{sim}(\mathbf{h}_{\mathbf{y}},\mathbf{h}_{\tilde{\mathbf{x}}})\,,$ (4)
the similarity (Eq. (2)) between their embeddings. The described procedure is depicted in Figure 1.
3.3 CompoNet
In order to showcase the utility of the COCOLA score as a metric for measuring the coherence of generated accompaniments, we propose an improved compositional model for music called CompoNet and compare it with the MSDM baseline in Section 5.
CompoNet first pre-trains a latent diffusion model $\epsilon_{\phi}$, based on U-Net architecture [42] conditioned via cross-attention layers [23], on a large dataset of tuples $(\mathbf{m},\mathbf{c})$ comprising audio mixtures $\mathbf{m}$ and relative textual descriptions $\mathbf{c}$. The mixtures $\mathbf{m}$ are mapped to latent vectors $\mathbf{z}=E^{\text{VAE}}(\mathbf{m})$ using a pre-trained VAE encoder [47], while the text descriptions are mapped to a continuous sequence $\mathbf{s}=E^{\text{TXT}}(\mathbf{c})$ using a text encoder (e.g., [10]). Following the DDPM formulation [7], the model is trained to reverse the forward Gaussian noising process given by:
$\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})\,,$ (5)
where $t\in[0,T]$ is a time index (with $T$ the maximum time step), $\bar{\alpha}_{t}$ is defined by integrating a noise schedule [7], and $\boldsymbol{\epsilon}$ is sampled from the standard Gaussian distribution. The model is trained with denoising score matching [39]:
$\min_{\phi}\mathbb{E}_{\mathbf{z},\boldsymbol{\epsilon},\mathbf{s},t}\left[\|\boldsymbol{\epsilon}-\epsilon_{\phi}(\mathbf{z}_{t},\mathbf{s},t)\|^{2}_{2}\right]\,,$ (6)
with $\mathbf{z}_{t}$ obtained via Eq. (5).
In a second phase, we fine-tune a ControlNet [17] adapter $C_{\psi}$ for tackling all compositional musical tasks (Figure 3) with a single model. The U-Net $\epsilon_{\phi}$ comprises an encoder, bottleneck, decoder structure $\epsilon_{\phi}=D_{\phi_{D}}\circ B_{\phi_{B}}\circ E_{\phi_{E}}$. The ControlNet adapter is defined as $C_{\psi}(\mathbf{z}_{t},\mathbf{s},t,\mathbf{w})=E_{\phi_{E}}(\mathbf{z}_{t}+\text{conv}_{\text{in}}(\mathbf{w}),\mathbf{s},t)$, where $\text{conv}_{\text{in}}$ is a zero-initialized convolutional layer, and $\mathbf{w}$ is a latent (VAE) embedding of an external audio input. The ControlNet adapter outputs the set of processed features $\{C^{i}_{\psi}(\mathbf{z}_{t},\mathbf{s},t,\mathbf{w})\}_{i=1,\dots,I}$ for each layer of $E_{\phi_{E}}$, with $I$ the total number of layers. The full ControlNet conditional architecture is defined as $\epsilon_{\phi,\psi}=D_{\phi_{D}}\circ B_{\phi_{B}}\circ E_{\phi_{E},\psi}$ with $E_{\phi_{E},\psi}$ combining the encoder and ControlNet adapter features at each layer $i$ with zero-initialized convolutional layers $\text{conv}_{i}$:
$E^{i}_{\phi_{E},\psi}(\mathbf{z}_{t},\mathbf{s},t,\mathbf{w})=E^{i}_{\phi_{E}}(\mathbf{z}_{t},\mathbf{s},t)+\text{conv}_{i}(C^{i}_{\psi}(\mathbf{z}_{t},\mathbf{s},t,\mathbf{w}))\,.$
While we have described the general ControlNet architecture, we still have to describe how we train it in CompoNet, namely, the roles of the $\mathbf{z},\mathbf{w}$ and $\mathbf{s}$ variables. Iterating over a dataset containing tuples $(\mathbf{x},\mathbf{t})$ with multi-stem tracks $\mathbf{x}=\{\mathbf{x}_{n}\}_{n=1,\dots,N}$ and tag descriptions $\mathbf{t}=\{\mathbf{t}_{n}\}_{n=1,\dots,N}$ for each stem, we sample from $\mathbf{x}$ two arbitrary subsets of stems $Y,X\subseteq\mathbf{x}$, with $|X|>0$. $Y$ contains input stems while $X$ contains output stems. The topological relationships between such subsets define all possible compositional tasks, as depicted in Figure 3. While previous models partially solve some tasks (see Table 1), ours is the first to solve all of them simultaneously. We proceed like in Eq. (1) and mix the sources in $Y$ and $X$, obtaining $\mathbf{m}_{Y}$ and $\mathbf{m}_{X}$, respectively. Afterward, we encode them in the VAE latent space, defining $\mathbf{z}=E^{\text{VAE}}(\mathbf{m}_{X})$ and $\mathbf{w}=E^{\text{VAE}}(\mathbf{m}_{Y})$. We define the following prompt $\mathbf{s}$:
$\mathbf{s}=E^{\text{TXT}}(\mathbf{t}_{Y_{1}},\dots,\mathbf{t}_{Y_{|Y|}},\verb|SEP|,\ \mathbf{t}_{X_{1}},\dots,\mathbf{t}_{X_{|X|}})\,,$ (7)
specifying input and output mixture tags separated by a special token SEP.
Having specified the required inputs and outputs, we train $\epsilon_{\phi,\psi}$ via Eq. (6), optimizing only the $\psi$ weights. The $\mathbf{s}$ prompt instructs the model which task to perform based on the specified stem tags. The user knows such information during the composition process in a generative DAW [48]. While we only test CompoNet on accompaniment generation against MSDM using COCOLA, we want the models to share similar abilities for a fair comparison.
4 Experimental Setup
Test Dataset
Train Dataset MUSDB18-HQ MoisesDB Slakh2100 CocoChorales
@@ 4 Experimental Setup
Table 2: Classification accuracy tests (%) with COCOLA models using $K=2$ sub-mixture test pairs (higher is better). MUSDB18-HQ is used as a hold-out test dataset.
4.1 Datasets
In our experiments, we use four different stem-separated public datasets for training COCOLA and fine-tuning CompoNet. The datasets are MUSDB18-HQ [52], MoisesDB [49], Slakh2100 [50] and CocoChorales [51].
MUSDB18-HQ [52] is the uncompressed version (in WAV format) of the MUSDB18 dataset, initially introduced in [53]. This dataset is a standard for evaluating music source separation systems. It comprises 150 tracks – 100 for training and 50 for testing – totaling approximately 10 hours of professional-quality audio. Each track is divided into four stems: Bass, Drums, Vocals, and Other, with “Other” covering any components not classified under the first three categories.
MoisesDB [49] features 240 music tracks across diverse genres and artists, accumulating more than 14 hours of music. Unlike MUSDB18-HQ, MoisesDB is a genuine multi-track dataset, offering a two-tier taxonomy of 11 distinct stems. Each stem in this dataset includes more detailed type annotations (e.g., Guitar might be labeled as Acoustic Guitar or Electric Guitar). Not having pre-computed splits, we set a custom 0.8 (train) / 0.1 (validation) / 0.1 (test) split.
Slakh2100 [50] is synthesized from the Lakh MIDI Dataset v0.1 [54] employing high-quality sample-based virtual instruments. It features 2100 tracks organized into 1500 tracks for training, 375 for validation, and 225 for testing, together amounting to 145h of audio. The tracks are annotated into 34 stem categories. While such a dataset contains an order of magnitude more data than MUSDB18-HQ and MoisesDB, it does not share the same level of realism as the latter, being the tracks synthesized from MIDI.
CocoChorales [51] is a chorale audio music dataset created through a synthesis process like Slakh2100. However, it comprises a substantially vaster collection of 240000 tracks, extending over 1411 hours of audio data. It is produced by generating symbolic notes via a Coconet model, performing their synthesis with MIDI-DDSP [55]. This dataset is richly annotated, featuring details on performance attributes and synthesis parameters. CocoChorales includes a diverse range of 13 instruments spanning Strings, Brass, Woodwind, and Random ensembles. In our experiments we use the tiny version, comprising 4000 tracks.
4.2 Model Implementation
To implement the COCOLA encoder $f_{\theta}$, we follow [33] and employ the EfficientNet-B0 [56] convolutional architecture followed by a linear projection layer, operating on the mel-filterbank audio representation. The embedding dimension is 512. With respect to the original baseline, we add a $0.1$ dropout on the EfficientNet layers.
For CompoNet, we employ the AudioLDM2 [57, 58] architecture. Since the authors pre-train the model on a large array of datasets [59, 60, 61, 62, 63, 64, 65, 66], we skip the pre-training phase in Section 3.2 and directly fine-tune a ControlNet adapter based on the AudioLDM 2-Large checkpoint [^3] https://huggingface.co/cvssp/audioldm2-large. During fine-tuning, we directly pass the conditioning prompt in Eq. 7 to the text-embedding mechanism of AudioLDM2 (based on CLAP [9, 27], T5 [10], and GPT2 [67]), conditioning both the U-Net and the ControlNet adapter.
100100100100200200200200300300300300400400400400500500500500600600600600700700700700800800800800000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111Training stepTrain accuracyCosine similarityBilinear similarity
Figure 4: Cosine vs bilinear similarity during training. Training is performed with “COCOLA All”. The accuracy metric is defined in Eq. (8). We adopt the bilinear similarity in Eq. (2), given its improved performance.
@@ 4.2 Model Implementation
CocoChorales
Slakh2100
MSDM [11] 0.23 92.81 2.01 3.31 0.72
CompoNet 0.30 106.23 3.20 13.50 8.32
Random 0.064 51.44 0.16 0.069 1.06
@@
MUSDB18-HQ
MSDM [11] 0.29 148.09 2.36 11.61 3.37
CompoNet 0.37 130.04 2.14 11.94 7.21
Random 0.11 100.25 0.35 4.40 3.47
@@
Table 3: Comparison between MSDM and CompoNet.
4.3 Training Details

### v3

@@ Abstract
We present COCOLA (Coherence-Oriented Contrastive Learning for Audio), a contrastive learning method for musical audio representations that captures the harmonic and rhythmic coherence between samples. Our method operates at the level of the stems composing music tracks and can input features obtained via Harmonic-Percussive Separation (HPS). COCOLA allows the objective evaluation of generative models for music accompaniment generation, which are difficult to benchmark with established metrics. In this regard, we evaluate recent music accompaniment generation models, demonstrating the effectiveness of the proposed method. We release the model checkpoints trained on public datasets containing separate stems (MUSDB18-HQ, MoisesDB, Slakh2100, and CocoChorales).
Index Terms:
Contrastive learning, generative AI, music information retrieval, performance evaluation
I Introduction
Recently, there have been significant advances in music generation in the continuous domain [1, 2, 3, 4], thanks to the impressive development of generative models [5, 6]. In addition to producing high-quality tracks of increasing length [3], these models offer precise semantic control through textual conditioning [7, 8]. However, they are limited as tools for musical composition, since they output a final mix containing all stems. To overcome this, a new range of compositional generative models for music is emerging [9, 10, 11], where (i) the generative tasks are defined at the stem level and (ii) their usage is iterative/interactive. The most important application of these models is accompaniment generation, where, given multiple conditioning sources, the model is asked to output a coherent stem that should combine with them in a harmonically and rhythmically consistent way, as opposed to being dissonant. While pioneering models [12, 13] could generate waveform accompaniments, their output is limited to a single stem category, and as such they cannot be used iteratively (acting sequentially on the model’s outputs) in a composition process.
@@ I Introduction
Figure 1: Illustration of COCOLA Score. COCOLA is a contrastive model that estimates the coherence between instrumental tracks and generated accompaniments.
To this end, we propose and release [^1] https://github.com/gladia-research-group/cocola a novel contrastive model called COCOLA (Coherence-Oriented Contrastive Learning for Audio), which can evaluate the coherence between conditioning tracks and generated accompaniments (Figure 1). The model is trained by maximizing the agreement between disjoint sub-components of an audio window (sub-mixtures of stems) and minimizing it on sub-components belonging to different windows. With the model, we define a COCOLA Score as the similarity between conditioning tracks and accompaniments in the embedding space, which we use to evaluate accompaniment generation models.
After discussing related work in Section II, we introduce COCOLA in Section III. We describe the experimental setup in Section IV and present the results in Section V. We conclude the article in Section VI.
Figure 2: The COCOLA training procedure (single stem case). We first randomly crop windows of size $L$ from a batch of $K$ tracks (depicted on the left). As a second step, we randomly select two distinct stems in each window. For example, in the first window, we select $\mathbf{x}^{1}_{1}$ (Guitar) and $\mathbf{x}^{1}_{3}$ (Drums). Thus, we embed all selected stems with the COCOLA encoder $f_{\theta}$, obtaining latent representations. For example, we obtain $\mathbf{h}^{1}_{1}$ and $\mathbf{h}^{1}_{2}$ from the first window. Finally, we compute the contrastive loss (Eq. (3)) considering embeddings belonging to the same window as positive pairs and combinations of embeddings between different windows as negative pairs.
II Related Work
II-A Contrastive Methods for Audio
@@
Supervised contrastive learning methods often rely on cross-modal approaches, requiring labeled data alongside audio. Early works [17] employ the contrastive loss [18] for aligning audio with simple labels. MuLaP [19] was the first model to jointly represent audio and complex text using a joint transformer encoder with cross-attention layers. Recent models [20, 7, 21] have adopted separate text and audio encoders, allowing independent use of these branches during inference.
@@
Self-supervised representation learning methods [22, 23] create embeddings from structural information in audio data. In [24], positive examples for triplet loss are built using Gaussian noise, time/frequency translations, and time proximity, without ensuring coherence (e.g., mixing unrelated sounds). Following methods [25] adapted the contrastive loss. COLA [26] uses the simple criterion of sampling positive pairs only from the same audio track, outperforming a supervised baseline. [27] pairs mixtures with sources extracted via source separation.
The proposed COCOLA method combines elements of both supervised and self-supervised approaches. Since stems are pre-separated, it is not fully self-supervised, but like self-supervised methods, it processes data using a uni-modal encoder.
II-B Waveform Music Accompaniment Generation Models
@@
MSDM [9], based on diffusion models [6, 5] was the first model able to generate waveform accompaniments belonging to different stem classes. Following, state-of-the-art models [11, 28] took the form of conditional generative models which respond with an output stem to an input track. Based on generative source separation via Bayesian inference [29, 30, 31], GMSDI [32] performs the tasks of MSDM, requiring models trained only with mixtures and text, by separating sources while generating them.
Another category of models [10, 33], following AUDIT [34], perform music editing by modifying the input track. Such models can add instruments to a track, but given the generative nature of the model, isolating such stems via difference with the input or source separation can introduce artefacts.
III Method
III-A Stem-Level Contrastive Learning
@@
$\mathbf{m}^{k}_{1}=\sum_{\mathbf{x}_{n}^{k}\in X_{1}^{k}}\mathbf{x}^{k}_{n},\qquad\mathbf{m}^{k}_{2}=\sum_{\mathbf{x}_{n}^{k}\in X_{2}^{k}}\mathbf{x}^{k}_{n}\,.$ (1)
When $X_{1}^{k},X_{2}^{k}$ are singletons, the sub-mixes are simply two stems in the window (single stem case).
Like in COLA[26], we use a convolutional audio-only encoder [^2] In our notation, we incorporate into $f_{\theta}$ any domain transform preceding or following the convolutional network operations, like the (pre) mel-filterbank map and the (post) projection head $g$ in COLA. $f_{\theta}:\mathbb{R}^{L}\to\mathbb{R}^{d}$, mapping $\mathbf{m}^{k}_{1}$ and $\mathbf{m}^{k}_{2}$ to lower-dimensional embedding vectors $\mathbf{h}^{k}_{1}=f_{\theta}(\mathbf{m}^{k}_{1})$ and $\mathbf{h}^{k}_{2}=f_{\theta}(\mathbf{m}^{k}_{2})$, with $d$ the embedding dimension. The COCOLA training procedure maximizes the agreement between pairs $\mathbf{h}^{k}_{1},\mathbf{h}^{k}_{2}$ of sub-mixes embeddings in the same window. It decreases it for pairs $\mathbf{h}^{k}_{1},\mathbf{h}^{j}_{2}$ ($j\neq k$) of sub-mixes embeddings in different windows. As in COLA, we use a bilinear similarity metric:
@@
TABLE I: Classification accuracy tests (%) with COCOLA models using $K=2$ sub-mixture test pairs (higher is better). MUSDB18-HQ is used as a hold-out test dataset.
Test Dataset
Model MUSDB18-HQ MoisesDB Slakh2100 CocoChorales
@@
w\ HPS
All 93.87 93.67 94.27 99.68
TABLE II: Classification accuracy tests (%) with COCOLA HPS All varying the values of $K$.
Test Dataset
K MUSDB18-HQ MoisesDB Slakh2100 CocoChorales
8 78.02 73.68 79.72 98.27
16 70.32 62.17 72.62 96.67
64 54.33 34.04 59.35 90.67
TABLE III: Comparison between music accompaniment generation models.
@@
Harmonic
MoisesDB Test
Random 0.072 19.37 0.38 50.53 52.34 53.86
@@
SA ControlNet [4, 42] 0.15 170.21 2.59 55.11 55.89 57.34
Diff-A-Riff [28] 0.14 12.16 0.90 57.34 58.00 60.00
@@
MUSDB18-HQ Test
Random 0.087 6.04 0.49 49.49 51.27 51.32
[…]
SA ControlNet [4, 42] 0.12 12.14 0.66 56.41 57.48 58.82
Diff-A-Riff [28] 0.20 119.41 1.68 53.79 54.75 56.08
[…]
In addition to the standard mel-filterbank representation of the input (as defined in COLA), we propose a variant where the COCOLA encoder operates on a pair of mel spectrogram features obtained via harmonic-percussive separation (HPS) [38]. As shown in Section V, this factorized input not only improves performance but also lets us better interpret which factor (harmony or rhythm) is more coherent.
III-B COCOLA Score
Equipped with the encoder $f_{\theta}$, we can quantify the coherence of the accompaniments generated by a generative model $p_{\phi}(\mathbf{x}\mid\mathbf{y})$, where $\mathbf{y}$ is the conditioning variable (the input) and $\mathbf{x}$ is the modeled variable (the output). The model’s variables can be either a set of stems or sub-mixes. Given the input $\mathbf{y}$, the model $p_{\phi}$ generates an output $\tilde{\mathbf{x}}\sim p_{\phi}(\mathbf{x}\mid\mathbf{y})$. We can compute the coherence between $\mathbf{y}$ and $\tilde{\mathbf{x}}$ by first embedding the two vectors $\mathbf{h}_{\mathbf{y}}=f_{\theta}(\mathbf{y})$ and $\mathbf{h}_{\tilde{\mathbf{x}}}=f_{\theta}(\tilde{\mathbf{x}})$ (summing the stems beforehand if considering a set of stems). We define the COCOLA Score between $\mathbf{x}$ and $\tilde{\mathbf{y}}$ as $\text{sim}(\mathbf{h}_{\mathbf{y}},\mathbf{h}_{\tilde{\mathbf{x}}})$, the similarity (Eq. (2)) between their embeddings. The described procedure is depicted in Figure 1.
Figure 3: Correlation plot between subjective scores (MOS) ($x$-axis) and COCOLA Scores ($y$-axis).
IV Experimental Setup
IV-A Datasets
In our experiments, we use four different stem-separated public datasets for training COCOLA. The datasets are MUSDB18-HQ [39], MoisesDB [35], Slakh2100 [36] and CocoChorales [37]. For MoisesDB, not having pre-determined splits, we set custom 0.8 (train) / 0.1 (validation) / 0.1 (test) splits. For CocoChorales, we use the tiny version, comprising a subset of 4000 tracks.
IV-B Model Implementation
To implement the COCOLA encoder $f_{\theta}$, we follow the COLA framework [26] and employ the EfficientNet-B0 [40] (using two input channels for HPS) convolutional architecture followed by a linear projection layer. The embedding dimension is 512. Differently from the original baseline, we add a $0.1$ dropout on the EfficientNet layers.
IV-C Training Details